# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

### v1

@@ Abstract
Encoding symmetries is a powerful inductive bias for improving the generalization of deep neural networks. However, most existing equivariant models are limited to simple symmetries like rotations, failing to address the broader class of general linear transformations, $\mathrm{GL}(n)$, that appear in many scientific domains. We introduce Reductive Lie Neurons (ReLNs), a novel neural network architecture exactly equivariant to these general linear symmetries. ReLNs are designed to operate directly on a wide range of structured inputs, including general $n$-by-$n$ matrices. ReLNs introduce a novel adjoint-invariant bilinear layer to achieve stable equivariance for both Lie-algebraic features and matrix-valued inputs, without requiring redesign for each subgroup. This architecture overcomes the limitations of prior equivariant networks that only apply to compact groups or simple vector data. We validate ReLNs’ versatility across a spectrum of tasks: they outperform existing methods on algebraic benchmarks with $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$ symmetries and achieve competitive results on a Lorentz-equivariant particle physics task. In 3D drone state estimation with geometric uncertainty, ReLNs jointly process velocities and covariances, yielding significant improvements in trajectory accuracy. ReLNs provide a practical and general framework for learning with broad linear group symmetries on Lie algebras and matrix-valued data. Project page: https://reductive-lie-neuron.github.io/
@@ 1 Introduction
Leveraging the symmetries present in real-world data improves the generalization of neural networks. When a model is equivariant, i.e., its outputs transform predictably with inputs under certain group actions, it can learn more efficiently and generalize across scenarios with less data.
Most advances in equivariant deep learning target compact symmetry groups, such as rotations, $\mathrm{SO}(3)$, or isometries, $\mathrm{E}(n)$, where strong geometric and algebraic tools support robust network design (Bronstein et al., 2021; Cohen & Welling, 2016; Thomas et al., 2018; Satorras et al., 2021; Finzi et al., 2020; Geiger & Smidt, 2022). These methods have led to practical gains. However, many real-world phenomena, such as affine distortions, sensor anisotropy, and general changes of basis, are governed by the broader general linear group $\mathrm{GL}(n)=\{A\in\mathbb{R}^{n\times n}:\det(A)\neq 0\}$, for which existing methods fail to provide exact equivariance. These more general symmetries are crucial in a wide array of scientific domains, including robotics, particle physics, and computer vision, a landscape we survey in Table 1.
Naive solutions, such as flattening matrices into vectors, discard essential geometric structure, while methods based on spectral decompositions are brittle and numerically unstable  (Magnus, 1985). This leaves a gap for a numerically stable neural architecture capable of processing diverse geometric data types equivariantly under arbitrary invertible linear transformations (Bronstein et al., 2021).
Figure 1: Examples of Lie groups and related manifolds in scientific applications. From left: the special linear group $\mathrm{SL}(3)$ (image homography), the Lorentz group $\mathrm{SO}(1,3)$ (spacetime symmetry), symplectic groups $\mathrm{Sp}(n)$ (Hamiltonian mechanics), the $\mathrm{SPD}(3)\oplus\mathbb{R}^{3}$ state space (probabilistic estimation), and the general linear group $\mathrm{GL}(3)$ (modeling stress-strain in continuum mechanics). See Table 1 for a more detailed survey.
Table 1: A survey of common Lie groups and their applications in equivariant deep learning.
Group / Data Structure Application Reference
$\mathrm{SO}(3)$ 3D Point Clouds, State Estimation Lin et al. (2024a); Deng et al. (2021); Son et al. (2024)
$\mathrm{SO}^{+}(1,3)$ Particle Physics, Jet Tagging Bogatskiy et al. (2020); Finzi et al. (2021); Batatia et al. (2023)
$\mathrm{SU}(3)$ Quantum Chromodynamics (QCD) Favoni et al. (2022)
$\mathrm{SL}(3)$ Homography Classification, 3D Vision Lin et al. (2024a); Finzi et al. (2021)
$\mathrm{Sp}(4,\mathbb{R})$ Hamiltonian dynamics Lin et al. (2024a); Finzi et al. (2021)
$\mathrm{SPD}(n)$ Geometric Uncertainty Processing -
$\mathrm{GL}(n)$ General Linear Transformations Basu et al. (2025); Finzi et al. (2021)
Not a group. $\mathrm{SPD}(n)$ is the manifold of symmetric positive-definite matrices, representable as the quotient space $\mathrm{GL}(n)/\mathrm{O}(n)$.
This work introduces Reductive Lie Neurons (ReLNs), a unified neural architecture that provides exact equivariance to the adjoint action of the general linear group, $\mathrm{GL}(n)$, and its reductive subgroups, enabling robust learning on both Lie-algebraic features and structured geometric data like covariances. We resolve the central technical obstacle for non-semisimple algebras—the degeneracy of the canonical Killing form—by introducing a learnable, provably non-degenerate, and $\mathrm{Ad}$-invariant bilinear form. This provides a mathematically stable mechanism for defining symmetry-preserving nonlinearities, pooling, and invariant layers applicable to the reductive Lie algebra $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$.
@@
    We propose Reductive Lie Neurons (ReLNs), a novel, general-purpose, and numerically stable network architecture for exact $\mathrm{GL}(n)$ adjoint equivariance.
@@
    We establish a connection between classical left-action equivariance and our adjoint-action framework for orthogonal groups. Through a provably-equivariant embedding map, we show that problems defined on standard vector actions—such as Lorentz transformations in particle physics or 3D point cloud processing—can be solved within our unified architecture, obviating the need for specialized model designs.
@@
    We establish the framework for geometric uncertainty-aware equivariant learning, enabling models to treat matrix-valued data that transforms under congruence (e.g., covariance tensors) as geometric objects.
  4. 4.
    We demonstrate the effectiveness of ReLNs through extensive experiments, showing that they outperform prior methods on Lie-algebraic benchmarks and achieve significant improvements in accuracy and robustness on a challenging 3D drone state estimation task.
@@ 2 Related Work
Encoding symmetry into neural architectures is a powerful inductive bias that improves data efficiency and generalization (Bronstein et al., 2021). In geometric deep learning, the most mature area of research focuses on enforcing equivariance to Euclidean isometries, transformations such as rotations and rigid motions formalized by Lie groups like $\mathrm{SO}(n)$ and $\mathrm{SE}(n)$. For grid-like data, foundational works include Group-Equivariant CNNs (Cohen & Welling, 2016) and Steerable CNNs (Weiler et al., 2018; Weiler & Cesa, 2021). For unstructured data, a dominant paradigm uses features associated with irreducible group representations (tensorial methods, E(n)-GNNs, and transformers (Thomas et al., 2018; Fuchs et al., 2020; Satorras et al., 2021; Batatia et al., 2022; Battiloro et al., 2025; Liao & Smidt, 2023; Assaad et al., 2023; Hutchinson et al., 2021)), with lightweight vector-based alternatives also available (Deng et al., 2021; Son et al., 2024). Theoretical analyses have also explored the universal approximation capabilities of such invariant networks (Maron et al., 2019). Complementing these specialized layers are general, model-agnostic strategies like frame averaging and canonicalization (Puny et al., 2022; Lin et al., 2024b; Kaba et al., 2023; Panigrahi & Mondal, 2024). Despite their success, most practical implementations are centered on scalar and vector features. While frameworks such as Tensor Field Networks (Thomas et al., 2018) or EMLP (Finzi et al., 2021) can handle higher-order tensor representations, scalable architectural design for general matrix-valued quantities (e.g., covariances or inertia tensors transforming as $\Sigma\mapsto R\Sigma R^{\top}$) is still limited in practice.
A major frontier is extending equivariance to non-compact groups like the general linear group $\mathrm{GL}(n)$ and the Lorentz group $\mathrm{SO}^{+}(1,3)$. Much of the work on the Lorentz group, which is critical in particle physics, has focused on designing specialized networks for the standard left action on vectors (Bogatskiy et al., 2020; Finzi et al., 2021; Batatia et al., 2023; Zhdanov et al., 2024). In contrast, we demonstrate that this left-action problem can be addressed within our universal, adjoint-equivariant framework via an embedding map.
One line of work generalizes group convolution and kernel design, often leveraging Fourier analysis (Xu et al., 2022; Helwig et al., 2023). This includes general frameworks for constructing equivariant networks on arbitrary matrix groups (Basu et al., 2025), through matrix functions (Batatia et al., 2024), and on reductive Lie groups (Batatia et al., 2023). Other approaches define kernels in the Lie algebra (Finzi et al., 2020), use Lie group decompositions for integration (Mironenco & Forré, 2024), or adapt canonicalization using infinitesimal generators of the Lie algebra (Shumaylov et al., 2025).
For these non-compact groups in general, the theoretical machinery underpinning many equivariant models does not readily apply. One line of work generalizes group convolution and kernel design, leveraging tools from Fourier analysis (Xu et al., 2022; Helwig et al., 2023), matrix functions (Batatia et al., 2024), or operating directly on reductive Lie groups (Batatia et al., 2023). Other methods focus on the Lie algebra, either by defining kernels within the algebra itself (Finzi et al., 2020) or using Lie group decompositions for integration (Mironenco & Forré, 2024; Shumaylov et al., 2025). Another approach leverages differential geometry, using tools such as partial differential operators (He et al., 2022; Shen et al., 2020; Jenner & Weiler, 2022) and the algebra of differential invariants (Sangalli et al., 2022; Li et al., 2024) to construct equivariant layers. A different generalist approach, taken by methods like the Equivariant MLP and G-RepsNet (Finzi et al., 2021; Basu et al., 2025), is to solve the equivariance constraint algebraically, though this often lacks specialized inductive biases like locality.
In contrast to these manifold-focused approaches, a parallel line of work operates directly on the Lie algebra. Lin et al. (2024a) introduced Lie Neurons, establishing a framework for adjoint-equivariance, but their method is restricted to semisimple Lie algebras where the Killing form is non-degenerate. This limitation precludes direct application to the reductive but non-semisimple algebra $\mathfrak{gl}(n)$, whose degenerate Killing form poses a central challenge. The high computational cost of group convolution methods, combined with the semisimple constraints of existing Lie-algebraic techniques, limits their direct applicability to robotics tasks that involve real-time processing, noisy measurements, or uncertainty estimation (Eschmann et al., 2024; Yu & Sun, 2024). While traditional, model-based algorithms like Kalman filters are designed to respect this geometry (Barrau & Bonnabel, 2016; Hartley et al., 2020), they lack the flexibility.
Our work confronts this challenge with ReLNs, a practical, numerically well-conditioned architecture for the full reductive group $\mathrm{GL}(n)$ and its subgroups. By introducing a learnable, non-degenerate, and $\mathrm{Ad}$-invariant bilinear form, our framework overcomes the key obstacle of the degenerate Killing form in $\mathfrak{gl}(n)$. This unified algebraic approach extends beyond strict Lie algebra elements to other geometric features, such as covariance matrices, that transform under congruence, providing a practical tool that sidesteps the complexities of prior methods requiring group integration or degenerate invariants. Figure 2 situates ReLNs as a general framework for $\mathrm{GL}(n)$ transformations, distinguishing our approach from architectures tailored to specialized subgroups.
@@ 3 Preliminaries
Our work builds equivariant networks on the Lie algebra $\mathfrak{g}=\mathrm{Lie}(G)$, the tangent space of a Lie group $G$ at the identity $e$. We focus on the general linear group $\mathrm{GL}(n)$ and its Lie algebra $\mathfrak{gl}(n)$, aiming for equivariance under the adjoint action $\mathrm{Ad}_{g}:X\mapsto gXg^{-1},\;g\in\mathrm{GL}(n),\;X\in\mathfrak{gl}(n)$. A central challenge arises from the structure of $\mathfrak{gl}(n)$: it is a reductive Lie algebra, and the canonical invariant inner product, the Killing form, is degenerate on $\mathfrak{gl}(n)$. This degeneracy poses a critical problem, as it prevents the construction of expressive nonlinear layers, causing them to collapse into linear maps and severely limiting the model’s expressive power. We address this by introducing a learnable, non-degenerate bilinear form, enabling fully nonlinear equivariant operations. For details of Lie theory and background, see Appendix A.
@@ 4 Reductive Lie Neurons: Architecture
We present ReLNs, a framework for building deep networks equivariant to the adjoint action of the general linear group $\mathrm{GL}(n)$. The design centers on a learnable, non-degenerate, $\mathrm{Ad}$-invariant bilinear form on the reductive Lie algebra $\mathfrak{gl}(n)$ and a complete toolbox of equivariant linear maps, nonlinearities, pooling, and invariant readouts.
4.1 A General $\mathrm{Ad}$-invariant Bilinear Form for Reductive Lie Algebras
The primary obstacle to applying Lie-algebraic methods such as Lie Neurons (Lin et al., 2024a) to $\mathfrak{gl}(n)$ is the degeneracy of its Killing form. We resolve this by constructing a modified bilinear form $\widetilde{B}$ that restores non-degeneracy while preserving the crucial $\mathrm{Ad}$-invariance property.

### v2

@@ Abstract
Many scientific and geometric problems exhibit general linear symmetries, yet most equivariant neural networks are built for compact groups or simple vector features, limiting their reuse on matrix-valued data such as covariances, inertias, or shape tensors. We introduce Reductive Lie Neurons (ReLNs), an exactly $\mathrm{GL}(n)$-equivariant architecture that natively supports matrix-valued and Lie-algebraic features. ReLNs resolve a central stability issue for reductive Lie algebras by introducing a non-degenerate adjoint (conjugation)-invariant bilinear form, enabling principled nonlinear interactions and invariant feature construction in a single architecture that transfers across subgroups without redesign. We demonstrate ReLNs on algebraic tasks with $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$ symmetries, Lorentz-equivariant particle physics, uncertainty-aware drone state estimation via joint velocity–covariance processing, learning from 3D Gaussian-splat representations, and EMLP double-pendulum benchmark spanning multiple symmetry groups. ReLNs consistently match or outperform strong equivariant and self-supervised baselines while using substantially fewer parameters and compute, improving the accuracy–efficiency trade-off and providing a practical, reusable backbone for learning with broad linear symmetries. Project page: https://reductive-lie-neuron.github.io/
@@ 1 Introduction
Exploiting symmetries in data is a fundamental principle of geometric deep learning. By enforcing equivariance, where model outputs transform predictably with inputs, neural networks can leverage strong inductive biases to improve data efficiency and generalization.
Substantial progress has been made for compact symmetry groups such as rotations $\mathrm{SO}(3)$ and Euclidean isometries $\mathrm{E}(n)$ (Bronstein et al., 2021; Cohen and Welling, 2016; Thomas et al., 2018; Satorras et al., 2021; Geiger and Smidt, 2022). However, many geometric and physical systems exhibit broader linear symmetries that are naturally expressed by the non-compact general linear group $\mathrm{GL}(n)=\{A\in\mathbb{R}^{n\times n}:\det(A)\neq 0\}$. Despite its ubiquity, general-purpose, exact equivariant architectures for $\mathrm{GL}(n)$ remain comparatively underdeveloped. Appendix D summarizes these groups and applications.
A key challenge for $\mathrm{GL}(n)$-equivariant learning is non-compactness. Unlike compact groups (e.g., $\mathrm{SO}(3)$), $\mathrm{GL}(n)$ does not come with a canonical $\mathrm{Ad}$-invariant positive-definite inner product: the Killing form is degenerate on $\mathfrak{gl}(n)$, and one loses the norm-preserving (unitary/orthogonal) structure that standard steerable irrep-based models rely on in practice (Thomas et al., 2018; Geiger and Smidt, 2022). At the Lie-algebra level, $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$ is reductive but not semisimple, so this degeneracy removes a canonical invariant metric and complicates stable invariant scalars for equivariant nonlinear. As a result, existing approaches either restrict attention to semisimple algebras or rely on ad hoc choices that do not provide a general, stable recipe for $\mathrm{GL}(n)$ adjoint equivariance.
Beyond this algebraic obstacle, real-world geometric inputs are often heterogeneous, with different attributes obeying different transformation laws. Vectors transform by a left action (e.g., $v\mapsto Av$), whereas matrix-valued tensors such as uncertainty, inertia, or shape covariances transform by congruence (e.g., $\Sigma\mapsto A\Sigma A^{\top}$). Learning to jointly couple vector–matrix channels in a symmetry-consistent and numerically stable way remains challenging: naive flattening discards tensor structure, and eigen/spectral parameterizations can be ambiguous and fragile under differentiation (Magnus, 1985). These issues arise in robotics state estimation, where velocities and covariance uncertainty must be processed together, and in 3D Gaussian Splatting, which couples a mean $\mu\in\mathbb{R}^{3}$ with an anisotropic covariance $\Sigma\in\mathrm{SPD}(3)$.
In this paper, we introduce Reductive Lie Neurons (ReLNs), a general-purpose architecture that is exactly equivariant to the adjoint action of $\mathrm{GL}(n)$ on $\mathfrak{gl}(n)$. ReLNs are built around a non-degenerate $\mathrm{Ad}$-invariant bilinear form that resolves Killing-form degeneracy on reductive algebras and enables stable invariant gating/normalization and expressive equivariant nonlinearities within a single backbone. We evaluate ReLNs on algebraic benchmarks ($\mathfrak{sl}(3)$, $\mathfrak{sp}(4)$), uncertainty-aware drone state estimation with joint velocity–covariance processing, learning from 3D Gaussian-splat representations with comparisons to Gaussian-MAE, and the EMLP double-pendulum benchmark; additional Lorentz-equivariant results are deferred to the appendix.
@@
    We propose Reductive Lie Neurons (ReLNs), a novel, general-purpose architecture for exact $\mathrm{GL}(n)$ adjoint equivariance, built upon a $\mathrm{Ad}$-invariant bilinear form that resolves the degeneracy issues of the standard Killing form on reductive algebras.
@@
    We introduce a unified lifting that embeds heterogeneous geometric inputs, left-acting vectors and congruence-transforming matrices (e.g., covariances, 3D Gaussian parameters), into a shared $\mathfrak{gl}(n)$ feature space, enabling a single adjoint-equivariant backbone with exact equivariance.
@@
    We demonstrate practical utility and efficiency across diverse regimes, matching Lie-algebra baselines on $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$, achieving strong robustness gains on uncertainty-aware drone state estimation and 3D Gaussian-splat representation learning, substantially reducing per-step compute compared to EMLP on the double-pendulum benchmark, matching baseline performance on jet top-tagging with Lorentz-equivariant embeddings.
Figure 1: Examples of Lie groups and related manifolds in scientific applications. From left: the special linear group $\mathrm{SL}(3)$ (image homography), the Lorentz group $\mathrm{SO}^{+}(1,3)$ (spacetime symmetry), symplectic groups $\mathrm{Sp}(n)$ (Hamiltonian mechanics), the $\mathrm{SPD}(3)\times\mathbb{R}^{3}$ state space (probabilistic estimation), and the general linear group $\mathrm{GL}(3)$ (modeling stress-strain in continuum mechanics).
@@ 2 Related Work
Encoding symmetry into neural architectures is a well-established inductive bias for improving data efficiency and generalization (Bronstein et al., 2021). Most geometric deep learning work targets Euclidean isometries such as rotations and rigid motions (e.g., $\mathrm{SO}(n)$ and $\mathrm{SE}(n)$). For grid data, representative approaches include group-equivariant and steerable CNNs (Cohen and Welling, 2016; Weiler et al., 2018; Weiler and Cesa, 2021). For sets and graphs, many methods build features from irreducible representations and tensorial message passing, including TFNs, $\mathrm{E}(n)$-GNNs, and equivariant transformers (Thomas et al., 2018; Fuchs et al., 2020; Satorras et al., 2021; Batatia et al., 2022; Liao and Smidt, 2023; Assaad et al., 2023; Hutchinson et al., 2021), alongside lighter vector-based variants (Deng et al., 2021; Son et al., 2024). Related theoretical work studies universality of invariant/equivariant architectures (Maron et al., 2019), and model-agnostic strategies such as canonicalization and frame averaging provide complementary routes to equivariance (Puny et al., 2022; Lin et al., 2024b; Kaba et al., 2023; Panigrahi and Mondal, 2024). While several frameworks can represent higher-order tensors (e.g., TFN (Thomas et al., 2018) and EMLP (Finzi et al., 2021)), scalable architectural design for general matrix-valued quantities (e.g., covariances transforming as $\Sigma\mapsto R\Sigma R^{\top}$) remains less standardized in practice.
Non-compact groups and $\mathrm{GL}(n)$. Extending equivariance to non-compact symmetries remains an active area, notably exemplified by Lorentz-equivariant architectures that operate on Minkowski geometry in particle physics (Bogatskiy et al., 2020; Gong et al., 2022; Batatia et al., 2023; Zhdanov et al., 2024). Complementary research utilizes geometric (Clifford) algebras to represent features as multivectors, leading to frameworks such as Geometric Algebra Transformers (GATr) Brehmer et al. (2023), Clifford-equivariant simplicial message passing Liu et al. (2024), and lightweight designs like GLGENN Filimoshina and Shirokov (2025). Approaches to non-compact groups include generalizations of group convolutions and kernels (Xu et al., 2022; Helwig et al., 2023), constructions based on matrix functions or reductive groups (Batatia et al., 2024; 2023), Lie-algebra-based kernels and decompositions (Finzi et al., 2020; Mironenco and Forré, 2024; Shumaylov et al., 2025), differential-operator and differential-invariant methods (He et al., 2022; Shen et al., 2020; Jenner and Weiler, 2022; Sangalli et al., 2022; Li et al., 2024), and algebraic constraint-solving frameworks such as EMLP and G-RepsNet (Finzi et al., 2021; Basu et al., 2025). These lines provide general tools but can incur substantial overhead (e.g., integration/basis computation) or lack inductive biases such as locality. Within Lie-algebraic adjoint-equivariant learning, Lie Neurons (Lin et al., 2024a) establish $\mathrm{Ad}$-equivariant neural network but rely on semisimple algebras where the Killing form is non-degenerate, and thus do not directly cover reductive algebras. This gap is relevant when inputs include general matrix-valued quantities that transform by adjoint action. Classical invariant filtering explicitly respects these transformation rules (Barrau and Bonnabel, 2016; Hartley et al., 2020), but is model-based and does not provide a learned backbone.
Learning on 3D Gaussian splats. 3D Gaussian Splatting represents geometry using anisotropic primitives parameterized by a mean $\mu\in\mathbb{R}^{3}$ and covariance $\Sigma\in\mathrm{SPD}(3)$. Recent learning frameworks over such primitives, including self-supervised learning and dynamics/segmentation models (Ma et al., 2025; Lu et al., 2024; Zhang et al., 2024; Ye et al., 2024), often treat heterogeneous attributes as loosely coupled channels (e.g., concatenating $\mu$ with rotation parameters and applying generic architectures), which does not explicitly enforce the coupled transformation of $(\mu,\Sigma)$. This decoupled treatment disregards the underlying geometry, representing vectors and tensors as mixed latent feature that fail to recognize their transformation laws. Consequently, such models struggle to capture the symmetry of the scene without massive data augmentation. ReLN formulation targets this gap by embedding these heterogeneous attributes into a unified Lie-algebraic space, enforcing the coupled equivariant structure of $(\mu,\Sigma)$, exploiting the isomorphism between congruence and adjoint actions under $\mathrm{SO(3)}$.
ReLNs overcome these limitations by establishing exact adjoint equivariance on $\mathfrak{gl}(n)$ via a non-degenerate $\mathrm{Ad}$-invariant bilinear form. This approach bypasses the reliance on degenerate invariants and group-specific integration required by previous methods. Figure 2 situates ReLNs relative to prior equivariant frameworks.
@@ 3 Preliminaries
We build equivariant networks on a reductive Lie algebra $\mathfrak{gl}(n)$, requiring equivariance under the adjoint action $\mathrm{Ad}_{g}(X)=gXg^{-1}$ for $g\in\mathrm{GL}(n)$. A key obstruction is that $\mathfrak{gl}(n)$ is reductive but not semisimple: its canonical $\mathrm{Ad}$-invariant symmetric bilinear form (the Killing form) is degenerate. Many equivariant architectures rely on a non-degenerate invariant form to produce invariant scalars (e.g., gating/normalization) and to parameterize genuinely nonlinear equivariant operations (e.g., Vector Neurons (Deng et al., 2021), Lie Neurons (Lin et al., 2024a)); degeneracy makes these constructions ill-conditioned or effectively linear. We resolve this by introducing a non-degenerate $\mathrm{Ad}$-invariant bilinear form on $\mathfrak{gl}(n)$, enabling fully nonlinear adjoint-equivariant layers. Our scope is reductive Lie algebras; extending to non-reductive cases (e.g., $\mathfrak{aff}(n)$) generally requires additional non-canonical completions. For details of Lie theory and background, see Appendix A.
@@ 4 Reductive Lie Neurons: Architecture
We present ReLNs, a framework for building deep networks equivariant to the adjoint action of $\mathrm{GL}(n)$ on its Lie algebra $\mathfrak{gl}(n)$. The design centers on a non-degenerate, $\mathrm{Ad}$-invariant bilinear form on the reductive Lie algebra $\mathfrak{gl}(n)$ and a complete toolbox of equivariant linear maps, nonlinearities, pooling, and invariant readouts.
4.1 An $\mathrm{Ad}$-invariant bilinear form for reductive Lie algebras
A key obstruction to extending Lie-algebraic designs (e.g., Lie Neurons (Lin et al., 2024a)) from semisimple Lie algebras to $\mathfrak{gl}(n)$ is that the Killing form may be degenerate on non-semisimple algebras. A finite-dimensional Lie algebra is reductive if it admits an ideal decomposition $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus[\mathfrak{g},\mathfrak{g}]$ with $\mathfrak{g},\mathfrak{g}]$ semisimple, and the Killing form is $B_{\mathfrak{g}}(X,Y)=\operatorname{Tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y})$; see Appendix A.4 for detailed definitions and the Killing form on $\mathfrak{g}$. In particular, $\mathfrak{gl}(n)$ is reductive but not semisimple: it contains a nontrivial center on which $\mathrm{ad}$ vanishes, forcing degeneracy of the Killing form. To obtain well-posed invariant contractions and nonlinearities, we construct an $\mathrm{Ad}$-invariant non-degenerate completion of the Killing form on any reductive Lie algebra.